For a fixed area, the length of a rectangle is inversly proportional to its width. A rectangle has a width of 18 feet and a length of 12 feet. If the length is increased to 16 feet, find the width.
Answer by RAY100(1637)

Length = K / Width
.
example is ,,,12 = K/18,,,,,K=216,,,
.
problem,,,,16=216/w,,,,,,w=13.5 ft
.
check: K=16*13.5 = 216,,,,ok
.